OpenStudy (anonymous):

Medal and fan Suppose you drop a tennis ball from a height of 6 feet. After the ball hits the floor, it rebounds to 80% of its previous height. How high will the ball rebound after its third bounce? Round to the nearest tenth.

OpenStudy (anonymous):

@Abdulhameed @Ethinzz Helpp?

OpenStudy (anonymous):

after first bounce 6*80/100 = 4.8 after second bounce 4.8*80/100= 3.84 after 3rd bounce 3.84*80/100= 3.072

OpenStudy (anonymous):

Thanks i appreciate it